OpenStudy (anonymous):

The scores of Tim’s first four math tests are 82%, 90%, 78%, and 100%. If Tim wants to have a test average of 90%, what must the average of Tim’s next two tests be? (A) 90% (B) 92% (C) 95% (D) 97% (E) 100%

OpenStudy (anonymous):

@UditKulka @nincompoop @kayla_b_stephens

OpenStudy (anonymous):

the average so far is (82 + 90 + 77 + 100) / 4

OpenStudy (anonymous):

87.5 % (i did use 78% for this result, typo above)

OpenStudy (anonymous):

so we need (87.5% + next_average ) / 2 =90%

OpenStudy (usarmy3947):

um @UditKulka can u please just give the link to the place ur taking the answer

OpenStudy (usarmy3947):

because i think steeling someone else's work is plagiarism

OpenStudy (usarmy3947):

here is the link :) http://openstudy.com/study#/updates/5246ece6e4b034054f3dd000

OpenStudy (anonymous):

90% is half of 180% Is this like 87.5%+92.5=180 180/2=90

OpenStudy (anonymous):

Will the answer be 92%?

OpenStudy (anonymous):

How many tests did he sit for?

OpenStudy (anonymous):

He took four tests so far.

OpenStudy (anonymous):

Each of them were out of 100 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So average in 4 tests is (82 + 90 + 77 + 100) / 4

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

I followed your method of adding 82%+90%+78%+100%=350% and then divided 350/4=87.5 However, I since he is taking another two tests after the four test, that means that it will be 6 tests altogether. To finding out the full percentage he needs I set up this equation of 90(6)6=540. Then subtracted 540-350=190. Since there will be two tests, I did 190/2=95%. I checked the answer key, it was 95%.

OpenStudy (anonymous):

yes sorry my bad

OpenStudy (anonymous):

No problem. I appreciate it when people at least try to help others.